# Sharp one-point estimates and Minkowski content for the scaling limit of three-dimensional loop-erased random walk

### v1

@@ 4 Sharp one-point function estimates
  • $S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset$.
The probability of these two events occurring is bounded above by $Cm^{-1}s^{\frac{L}{4}-1}$, which is, taking $L$ sufficiently large, equal to
$P^{y,y}_{1,2}\Big{(}\text{LE}\big{(}S^{1}[0,\infty)\big{)}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}\Big{[}1+O\big{(}s\big{)}\Big{]}.$
@@
$P^{y,y}_{1,2}\Big{(}\gamma^{1}[a,b]\cap S^{2}[t_{2,L},t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
$=P^{y,y}_{1,2}\Big{(}\gamma^{1}[a,b]\cap S^{2}[1,t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}\Big{[}1+O\big{(}s\big{)}\Big{]}.$
Now we can couple $\gamma^{1}$ and its scaling limit ${\eta}_{\infty}$. We also couple $S^{2}$ and the Brownian motion $W$. By the invariance under rotation, taking $m$ and $L$ sufficiently large, we see that
$P^{y,y}_{1,2}\Big{(}\gamma^{1}[a,b]\cap S^{2}[t_{2,L},t_{2}]=\emptyset,\ S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
$=\Big{[}1+O\big{(}s\big{)}\Big{]}\times\sum_{w\in\partial B\big{(}x_{0},s+s^{L}\big{)}\cap B(y,s^{L/2})}P^{y}\Big{(}S^{2}[1,t_{2,L}]\cap B(x_{0},s)=\emptyset,\ S^{2}(t_{2,L})=w\Big{)}$
×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times P^{y,w}_{1,2}\Big{(}\gamma^{1}[a,b]\cap S^{2}[0,t_{2}]=\emptyset,\ S^{2}[0,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
$=\Big{[}1+O\big{(}s\big{)}\Big{]}P^{y}\Big{(}S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
×$\ \ \ \ \ \ \ \ \ \ \ \times P^{y,w_{y}}_{1,2}\Big{(}\gamma^{1}[a,b]\cap S^{2}[0,t_{2}]=\emptyset\ \Big{|}\ S^{2}[0,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
$=\Big{[}1+O(s)\Big{]}P^{y}\Big{(}S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
×$\ \ \ \ \ \ \ \ \ \ \ \times P^{y_{0},w_{y_{0}}}_{1,2}\Big{(}\gamma^{1}[a,b]\cap S^{2}[0,t_{2}]=\emptyset\ \Big{|}\ S^{2}[0,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}.$
In the second equality, $w_{y}$ is the unique point in $\partial B\big{(}x_{0},s+s^{L}\big{)}\cap B(y,s^{L/2})$ such that $x_{0}$, $y$ and $w_{y}$ are lying on the same straight line. In the third equality, $y_{0}\in\partial_{i}B(x_{0},s)$ is the “north pole” of $B(x_{0},s)$.
Therefore, returning to (4.2), it follows that for all $y,z\in\partial_{i}B(x_{0},s)$
$\frac{P^{y,y}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}}{P^{z,z}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}}$
$=\Big{[}1+O\big{(}s^{c_{0}}\big{)}\Big{]}\frac{P^{y}\Big{(}S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}}{P^{z}\Big{(}S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}}.$
@@
$P\Big{(}\gamma_{m}\cap B(\hat{x},s)\neq\emptyset\Big{)}$
$=P^{y_{0},y_{0}}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}$
×$\times\Big{[}1+O\big{(}s^{c_{0}}\big{)}\Big{]}\sum_{y\in\partial_{i}B(x_{0},s)}\frac{P^{y}\Big{(}S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}}{P^{y_{0}}\Big{(}S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}}.$
@@
$\sum_{y\in\partial_{i}B(x_{0},s)}\frac{P^{y}\Big{(}S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}}{P^{y_{0}}\Big{(}S^{2}[1,t_{2}]\cap B(x_{0},s)=\emptyset\Big{)}}=c_{1}s^{2}m^{2}\Big{[}1+O\big{(}s^{c_{2}}\big{)}\Big{]}.$
@@
$P^{y_{0},y_{0}}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ \text{LE}\big{(}S^{1}[0,t^{1}_{0}]\big{)}[u_{1},u_{2}]\cap S^{2}[1,T^{2}]=\emptyset,\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}.$
But, it is not difficult to show that for sufficiently large $m$
$P^{y_{0},y_{0}}_{1,2}\Big{(}t^{1}_{0}<T^{1},\ S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset\Big{)}=q_{1}s^{-1}m^{-2}\Big{[}1+O\big{(}s^{q_{2}}\big{)}\Big{]},$
@@
where $X^{1}$ is $S^{1}$ conditioned that $t^{1}_{0}<T^{1}$ and $X^{2}$ is $S^{2}$ conditioned that $S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset$. We obtain such estimate in Lemma 4.8, and this completes the proof of the proposition. ∎
We finish this subsection with the following lemma for the random walks $S^{1}$ and $S^{2}$ conditioned to the events appearing in (4.26). Let $X^{1}$ be the simple random walk $S^{1}$ conditioned to $t^{1}_{0}<T^{1}$, i.e. $X^{1}$ hits the origin before $\mathbb{D}$. Let $X^{2}$ be the simple random walk $S^{2}$ conditioned that $S^{2}[1,T^{2}]\cap B(x_{0},s)=\emptyset$, then $X^{2}$ does not return to the ball $B(x_{0},s)$. We show next the sharp estimate on their intersection probability.
@@
Recall the definitions of $u_{2}$ and $u_{1}$ in (4.19) and (4.20), respectively. Let us write
@@
We first deal with the case that $s=2^{-k}$. The coupling argument shows that, for an infinite LERW in $m^{-1},\mathbb{Z}^{3}$ $\gamma^{1}=\text{LE}\big{(}S^{1}[0,\infty)\big{)}$,
$\frac{p(2^{-k-1})}{p(2^{-k})}=\frac{P^{y_{0},y_{0}}_{1,2}\Big{(}\gamma^{1}[a,b]\cap X^{2}[1,t_{2}]=\emptyset\Big{)}}{P^{y_{0}^{\prime},y_{0}^{\prime}}_{1,2}\Big{(}\gamma^{1}[a^{\prime},b]\cap X^{2}[1,t_{2}]=\emptyset\Big{)}}\ \Big{[}1+O\big{(}2^{-ck}\big{)}\Big{]},$
@@
  • $y_{0}$ is the north pole of $B(x_{0},2^{-k-1})$ and $y_{0}^{\prime}$ is the north pole of $B(x_{0},2^{-k})$
  • $b=\inf\Big{\{}j\geq 0\ \Big{|}\ \gamma^{1}(j)\notin B\big{(}x_{0},2^{-\delta_{0}k}\big{)}\Big{\}}$,
  • $a=\max\Big{\{}j\in[0,b]\ \Big{|}\ \gamma^{1}(j)\in B(x_{0},2^{-k-1})\Big{\}}$,
  • $a^{\prime}=\max\Big{\{}j\in[0,b]\ \Big{|}\ \gamma^{1}(j)\in B(x_{0},2^{-k})\Big{\}}$ and
  • $t_{2}=\inf\Big{\{}j\geq 0\ \Big{|}\ X^{2}(j)\notin B\big{(}x_{0},2^{-\delta_{0}k}\big{)}\Big{\}}$.
The same coupling argument gives that
$\frac{p(2^{-k})}{p(2^{-k+1})}=\frac{P^{y_{0}^{\prime},y_{0}^{\prime}}_{1,2}\Big{(}\gamma^{1}[a^{\prime\prime},b^{\prime}]\cap X^{2}[1,t_{2}^{\prime}]=\emptyset\Big{)}}{P^{y_{0}^{\prime\prime},y_{0}^{\prime\prime}}_{1,2}\Big{(}\gamma^{1}[a^{\prime\prime\prime},b^{\prime}]\cap X^{2}[1,t_{2}^{\prime}]=\emptyset\Big{)}}\ \Big{[}1+O\big{(}2^{-ck}\big{)}\Big{]},$
@@
  • $b^{\prime}=\inf\Big{\{}j\geq 0\ \Big{|}\ \gamma^{1}(j)\notin B\big{(}x_{0},2^{-\delta_{0}k-1}\big{)}\Big{\}}$,
  • $a^{\prime\prime}=\max\Big{\{}j\in[0,b^{\prime}]\ \Big{|}\ \gamma^{1}(j)\in B(x_{0},2^{-k})\Big{\}}$,
  • $a^{\prime\prime\prime}=\max\Big{\{}j\in[0,b]\ \Big{|}\ \gamma^{1}(j)\in B(x_{0},2^{-k+1})\Big{\}}$ and
  • $t_{2}^{\prime}=\inf\Big{\{}j\geq 0\ \Big{|}\ X^{2}(j)\notin B\big{(}x_{0},2^{-\delta_{0}k-1}\big{)}\Big{\}}$.
Replacing $\gamma^{1}$ and $X^{2}$ by their scaling limits which satisfy the scale invariance, we see that
@@
We can again use the scale invariance of the scaling limits of $\gamma^{1}$ and $X^{2}$ to show that
@@
This implies (4.30). Furthermore, it is not difficult to prove that $f$ is continuous. Combining these properties with $f(0)=1$ and $f(1)=2^{-\alpha}$. it holds that $f(r)=2^{-\alpha r}$. Consequently, we have
@@ 4.3 Asymptotics of the one-point function
By continuity and rotational invariance of the one-point fuction (Theorem 1.3), in order to prove (1.12) it suffices to verify that for some $b_{1}>0$
@@
From which (4.32) follows and hence (1.12).
Given 1.15, a similar argument holds for the asymptotic (1.13) as $|x|\rightarrow 1$. ∎
@@ 5.1 An $L^{2}$ estimate
Consider a fixed $V\in{\cal D}^{o}$ and let
$J_{s}(z)\coloneqq 2^{(3-\beta)s}1_{z\in B({\cal K},2^{-s})}$ (5.1)
be the renormalized indicator function of the blow-up of ${\cal K}$. Write
[…]
@@
$E\left[(J_{s}(V)-J_{s^{\prime}}(V))^{2}\right]=O_{V}\left(2^{-cs^{1/8}}\right).$ (5.2)
@@
[…]
@@
[…]

### v2

@@ 4 Sharp one-point function estimates
  • $S^{2}[1,t_{2}]\cap B(\hat{x},s)=\emptyset$.
By classical estimates on escape probabilities for SRW, the probability of these two events occurring is bounded above by $Cm^{-1}s^{\frac{L}{4}-1}$, which is, taking $L$ sufficiently large, equal to
$P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)O\big(s\big).$
@@
$P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}[t_{2,L},t_{2}]=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)$
$=P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}_{-}=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)\Big[1+O\big(s\big)\Big].$
Now we can couple $\gamma^{\infty}$ and its scaling limit $\eta$$_{\infty}$. We also couple $S^{2}$ and the Brownian motion $W$. Taking $m$ and $L$ sufficiently large, we see that
$P^{y,y}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}[t_{2,L},t_{2}]=\emptyset,\ S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)$
$=\Big[1+O\big(s\big)\Big]\times\sum_{w\in\partial B\big(\hat{x},s+s^{L}\big)\cap B(y,s^{L/2})}P^{y}\Big(S^{2}[1,t_{2,L}]\cap B(\hat{x},s)=\emptyset,\ S^{2}(t_{2,L})=w\Big)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times P^{y,w}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}[0,t_{2}]=\emptyset,\ S^{2}[0,t_{2}]\cap B(\hat{x},s)=\emptyset\Big)$
$=\Big[1+O\big(s\big)\Big]P^{y}\Big(S^{2}[1,t_{2}]\cap B(\hat{x},s)=\emptyset\Big)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times P^{y,w_{y}}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}[0,t_{2}]=\emptyset\ \Big|\ S^{2}[0,t_{2}]\cap B(\hat{x},s)=\emptyset\Big)$
$=\Big[1+O(s)\Big]P^{y}\Big(S^{2}[1,t_{2}]\cap B(\hat{x},s)=\emptyset\Big)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times P^{y_{0},w_{y_{0}}}_{1,2}\Big(\gamma^{\infty}[a,b]\cap S^{2}[0,t_{2}]=\emptyset\ \Big|\ S^{2}[0,t_{2}]\cap B(\hat{x},s)=\emptyset\Big),$
where in the second equality, $w_{y}$ is the unique point in $\partial B\big(\hat{x},s+s^{L}\big)\cap B(y,s^{L/2})$ such that $\hat{x}$, $y$ and $w_{y}$ are lying on the same straight line and in the third equality, $y_{0}\in\partial_{i}B(\hat{x},s)$ is the “north pole” of $B(\hat{x},s)$. Note that the third equality follows from approximation of non-intersection probabilities by the continuous counterpart (similar to the argument in [20, Proposition 3.16]) and the invariance under rotation of $\eta_{\infty}$ and $W$.
Therefore, returning to (4.2), it follows that for all $y,z\in\partial_{i}B(\hat{x},s)$,
$\begin{split}\frac{P^{y,y}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)}{P^{z,z}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)}\\ =\Big[1+O\big(s^{c_{0}}\big)\Big]\frac{P^{y}\Big(S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}{P^{z}\Big(S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}.\end{split}$
@@
$P\Big(\gamma_{m}\cap B(\hat{x},s)\neq\emptyset\Big)$ $=P^{y_{0},y_{0}}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)$
$\qquad\qquad\times\Big[1+O\big(s^{c_{0}}\big)\Big]\sum_{y\in\partial_{i}B(\hat{x},s)}\frac{P^{y}\Big(S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}{P^{y_{0}}\Big(S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}.$
@@
$\sum_{y\in\partial_{i}B(\hat{x},s)}\frac{P^{y}\Big(S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}{P^{y_{0}}\Big(S^{2}_{-}\cap B(\hat{x},s)=\emptyset\Big)}=c_{1}s^{2}m^{2}\Big[1+O\big(s^{c_{2}}\big)\Big].$
@@
$P^{y_{0},y_{0}}_{1,2}\Big(t^{1}_{0}<T^{1},\ \gamma^{0}[u_{1},u_{2}]\cap S^{2}_{+}=\emptyset,\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big).$
But, by classical hitting probability estimates and asymptotics for Green’s function, it is not difficult to show that for sufficiently large $m$
$P^{y_{0},y_{0}}_{1,2}\Big(t^{1}_{0}<T^{1},\ S^{2}_{+}\cap B(\hat{x},s)=\emptyset\Big)=q_{1}s^{-1}m^{-2}\Big[1+O\big(s^{q_{2}}\big)\Big],$
@@
where $X^{1}$ is $S^{1}$ conditioned that $t^{1}_{0}<T^{1}$ and $X^{2}$ is $S^{2}$ conditioned that $S^{2}_{+}\cap B(\hat{x},s)=\emptyset$. We obtain such estimate in Lemma 4.8, and this completes the proof of the proposition. ∎
We finish this subsection with the following lemma for the random walks $S^{1}$ and $S^{2}$ conditioned to the events appearing in (4.37). Let $X^{1}$ be the simple random walk $S^{1}$ conditioned to $t^{1}_{0}<T^{1}$, i.e. $X^{1}$ hits the origin before $\mathbb{D}$. Let $X^{2}$ be the simple random walk $S^{2}$ conditioned that $S^{2}_{+}\cap B(\hat{x},s)=\emptyset$, then $X^{2}$ does not return to the ball $B(\hat{x},s)$. We show next the sharp estimate on their intersection probability.
@@
Recall the definitions of and $u_{1}$ and $u_{2}$ in (4.23) and (4.24), respectively. Let us write
@@
We first deal with the case that $s=2^{-k}$. Comparison results between the loop-erasure of $X^{1}$ and an infinite LERW (see Propositions 4.2, 4.4 and Corollary 4.5 of [23] for this) show that, for an infinite LERW in $m^{-1}\mathbb{Z}^{3}$, again denoted by $\gamma^{\infty}=\text{LE}\big(S^{1}[0,\infty)\big)$,
$\frac{p(2^{-k-1})}{p(2^{-k})}=\frac{P^{y_{0},y_{0}}_{1,2}\Big(\gamma^{\infty}[a,b]\cap X^{2}[1,t_{2}]=\emptyset\Big)}{P^{y_{0}^{\prime},y_{0}^{\prime}}_{1,2}\Big(\gamma^{\infty}[a^{\prime},b]\cap X^{2}[1,t_{2}]=\emptyset\Big)}\ \Big[1+O\big(2^{-ck}\big)\Big],$
@@
  • $y_{0}$ is the north pole of $B(\hat{x},2^{-k-1})$ and $y_{0}^{\prime}$ is the north pole of $B(\hat{x},2^{-k})$
  • $b=\inf\Big\{j\geq 0\ \Big|\ \gamma^{\infty}(j)\notin B\big(\hat{x},2^{-\delta_{0}k}\big)\Big\}$,
  • $a=\max\Big\{j\in[0,b]\ \Big|\ \gamma^{\infty}(j)\in B(\hat{x},2^{-k-1})\Big\}$,
  • $a^{\prime}=\max\Big\{j\in[0,b]\ \Big|\ \gamma^{\infty}(j)\in B(\hat{x},2^{-k})\Big\}$ and
  • $t_{2}=\inf\Big\{j\geq 0\ \Big|\ X^{2}(j)\notin B\big(\hat{x},2^{-\delta_{0}k}\big)\Big\}$.
The same comparison argument gives that
$\frac{p(2^{-k})}{p(2^{-k+1})}=\frac{P^{y_{0}^{\prime},y_{0}^{\prime}}_{1,2}\Big(\gamma^{\infty}[a^{\prime\prime},b^{\prime}]\cap X^{2}[1,t_{2}^{\prime}]=\emptyset\Big)}{P^{y_{0}^{\prime\prime},y_{0}^{\prime\prime}}_{1,2}\Big(\gamma^{\infty}[a^{\prime\prime\prime},b^{\prime}]\cap X^{2}[1,t_{2}^{\prime}]=\emptyset\Big)}\ \Big[1+O\big(2^{-ck}\big)\Big],$
@@
  • $b^{\prime}=\inf\Big\{j\geq 0\ \Big|\ \gamma^{\infty}(j)\notin B\big(\hat{x},2^{-\delta_{0}k-1}\big)\Big\}$,
  • $a^{\prime\prime}=\max\Big\{j\in[0,b^{\prime}]\ \Big|\ \gamma^{\infty}(j)\in B(\hat{x},2^{-k})\Big\}$,
  • $a^{\prime\prime\prime}=\max\Big\{j\in[0,b]\ \Big|\ \gamma^{\infty}(j)\in B(\hat{x},2^{-k+1})\Big\}$ and
  • $t_{2}^{\prime}=\inf\Big\{j\geq 0\ \Big|\ X^{2}(j)\notin B\big(\hat{x},2^{-\delta_{0}k-1}\big)\Big\}$.
Replacing $\gamma^{\infty}$ and $X^{2}$ by their scaling limits which satisfy the scale invariance, we see that
@@
We can again use the scale invariance of the scaling limits of $\gamma^{\infty}$ and $X^{2}$ to show that
@@
This implies (4.41). Furthermore, using the coupling argument as in Step 2 of the proof of Proposition 3.1, it is not difficult (but rather tedious) to prove that $f$ is continuous. Combining these properties with $f(0)=1$ and $f(1)=2^{-\alpha}$. it holds that $f(r)=2^{-\alpha r}$. Consequently, we have
@@ 4.3 Asymptotics of the one-point function
By continuity and rotational invariance of the one-point function (Theorem 1.3), in order to prove (1.12) it suffices to verify that for some $b_{1}>0$
@@
On the other hand, a similar argument used to prove (4.12) gives that for each $r\in[0,1]$
$\frac{g\big(2^{-r}\cdot 2^{-n}\hat{x}\big)}{g(2^{-n}\hat{x})}=2^{r(3-\beta)}\Big[1+O(2^{-\delta n})\Big],\ \ \ \ \text{as}\ \ n\to\infty.$
Combining these estimates, we obtain (4.43) and hence (1.12).
Given (1.11), a similar argument holds for the asymptotic (1.13) as $|x|\rightarrow 1$. ∎
@@ 5.1 An $L^{2}$ estimate
Consider a fixed $V\in{\cal D}^{o}$ and recall Proposition 2.2 (with $A$, $\delta$, $d$ therein as $\cal K$}, $\beta$, $3$ resp.) for
$J_{s}(z)\coloneqq 2^{(3-\beta)s}1_{z\in B({\cal K},2^{-s})},$ (5.1)
the renormalized indicator function of the blow-up of $\cal K$}. Write
$J_{s}(V)=\int_{V}J_{s}(z)dz$
@@
$E\left[(J_{s}(V)-J_{s^{\prime}}(V))^{2}\right]=O_{V}\left(2^{-cs^{1/2}}\right).$ (5.2)
@@
$\begin{split}E\left[(J_{s}(V)-J_{s^{\prime}}(V))^{2}\right]=\;&\iint_{z,w\in V,|z-w|\geq 2^{-\sqrt{s}/100}}E[K]dzdw\\ +\;&\iint_{z,w\in V,100\cdot 2^{-s}\leq|z-w|\leq 2^{-\sqrt{s}/100}}E[K]dzdw\\ +\;&\iint_{z,w\in V,|z-w|\leq 100\cdot 2^{-s}}E[K]dzdw\\ \coloneqq\;&{\rm I}+{\rm II}+{\rm III}.\end{split}$ (5.4)
@@
$\begin{split}\big|{\rm II}\big|\;&\leq\iint_{z,w\in V,100\cdot 2^{-s}\leq|z-w|\leq 2^{-\sqrt{s}/100}}E\Big[\big(J_{s}(z)+J_{s^{\prime}}(z)\big)\big(J_{s}(w)+J_{s^{\prime}}(w)\big)\Big]dzdw\\ &\leq c\int_{V}\int_{100\cdot 2^{-s}}^{2^{-\sqrt{s}}}\int_{\partial B(z,r)}\left(2^{(3-\beta)s}\right)^{2}\cdot\left(\frac{2^{-2s}}{r}\right)^{3-\beta}\sigma_{r}(dw)drdz\\ &\leq C_{V}2^{-(\beta-1)\sqrt{s}/100},\end{split}$ (5.5)